OpenStudy (anonymous):

using the zero principle (9r+4)(5r-15)=0

OpenStudy (anonymous):

just equate the 2 factors by 0.

OpenStudy (anonymous):

making 9r + 4 = 0 and 5r - 15 = 0

OpenStudy (anonymous):

Solve it on your own.

OpenStudy (anonymous):

would it be -4,9 and 5,15 or is it a fraction.

OpenStudy (anonymous):

the first equation would be a fraction but the 2nd would be an integer

OpenStudy (anonymous):

so it would be -4/9, 15 I am lost I think.

OpenStudy (anonymous):

you got the -4/9, @ the 2nd equation, try to transpose -15 to the other side and divide by 5 to get the answer.

OpenStudy (anonymous):

so would this be the same principle t^2+8t+15=0

OpenStudy (anonymous):

this is a different equation

OpenStudy (anonymous):

yes. Factor the quadratic equation and then equate both with 0 then solve. :)